7. The data below are the final exam scores of 10 randomly selected statistics students and the number of hours they studied for the exam. Find the standard error of estimate, se, given that Hours (X) 3 5 2 8 2 4 4 5 6 3

Scores (Y) 65 80 60 88 66 78 85 90 90 71

A. 6.305

B. 9.875

C. 8.912

D. 7.913
Answer #1

From the given data, we get the regression equation as below (The same is calculated after the answer.

y' = 56.1139 + 5.0443x

Standard error of the estimate = SQRT {SUM[(y - y')2] / N - 2}. Here N = 10

Calculation of SUM(y - y')2. We put each value of x in the regression equation to get y'.

x y y' y - y' (y - y')2
3 65 71.2468 6.2468 39.02251
5 80 81.3354 1.3354 1.783293
2 60 66.2025 6.2025 38.47101
8 88 96.4683 8.4683 71.7121
2 66 66.2025 0.2025 0.041006
4 78 76.2911 -1.7089 2.920339
4 85 76.2911 -8.7089 75.84494
5 90 81.3354 -8.6646 75.07529
6 90 86.3797 -3.6203 13.10657
3 71 71.2468 0.2468 0.06091
318.038

Therefore the Se = Sqrt(318.038 / 8) = 6.305 (Option A)

______________________________________________

The Regression Equation:

From the given data, we get:

Count x y xy x2 y2
1 3 65 195 9 4225
2 5 80 400 25 6400
3 2 60 120 4 3600
4 8 88 704 64 7744
5 2 66 132 4 4356
6 4 78 312 16 6084
7 4 85 340 16 7225
8 5 90 450 25 8100
9 6 90 540 36 8100
10 3 71 213 9 5041
Total 42 773 3406 208 60875

Sum x = 42, Sum y = 773, Sum xy = 3406, Sum x2 = 208, (Sum x)2 = 1764, n = 10

b1: The slope


b0: The coefficient


Therefore the regression equation = y' = b0 + b1x

y' = 56.1139 + 5.0443x